Which of the following are the factors of x^3 + 3x^2?
Found 2 solutions by checkley71, Nate:
Answer by checkley71(8403)

x~3+3x~2 or x~2(x+3) or x~2=0 or x=0 and x+3=0 or x=-3
PROOF 0~3+3(0)~2 = 0 and -3~3+3(-3)~2 = 0 OR -27 +3(-3)~2 = 0 OR -27 +3(9) =0
OR -27+27=0.

Answer by Nate(3500)

The only factors for the expression x^3 + 3x^2 is x^2 and x+3
x^3 + 3x^2 original
X^2(x+3)
Since this is not an equation problem, you do not have x equalling a number.